# PointDreamer: Zero-shot 3D Textured Mesh Reconstruction from Colored Point Cloud by 2D Inpainting

### v1

@@ 3.2.1. Projection: 3D to 2D
Directly projecting the whole point cloud into 2D would incorporate points from occluded parts that should not be visible from the given view. This would mess up the inpainting results and subsequent mesh textures; see the bottom of Figure 3 as an example. Therefore, before generating the sparse images, we precess the input point cloud by the “Hidden Point Removal” operator (Katz et al., 2007), which works by transforming the input and extracting the points that reside on its convex hull. With pre-set camera parameters of $K$ viewpoints ($K=8$ in our implementation) and associated visible point clouds, we conduct camera transformation to these points to get their corresponding pixel coordinates in 2D image space. These pixels are painted according to their associated 3D points’ colors, producing $K$ sparse images. Comparisons of different $K$ values can be found in our supplementary file.
3.2.2. Inpainting: Sparse to Dense
@@
Any 2D inpainting method that fills empty pixels in input images can serve as our inpainting module. We propose to use the state-or-the-art DDNM (Wang et al., 2022b) based on a pre-trained unconditional 2D diffusion model (Dhariwal and Nichol, 2021) to inpaint the multi-view sparse images into dense ones $\mathbf{I}=\{I_{k}\}_{k=1}^{K}$. In this way, the strong prior provided by the 2D diffusion model can facilitate high-quality inpainting. So far, we have completed color prediction purely in 2D space instead of 3D or UV space.
3.3. Unprojection: 2D to 3D
@@
The top of Figure 4 illustrates the basic idea of the unprojection module. As can be seen, we represent mesh texture as a texture atlas $\mathcal{T}$ by applying UV mapping to a mesh $\mathcal{M}$ via Xatlas (Young, 2022), following (Gao et al., 2022; Long et al., 2023). A texture atlas is an RGB image containing different charts. Each chart is a continuous area in UV space that corresponds to a continuous segment of $\mathcal{M}$ in 3D space. Each pixel $\mathcal{T}[u,v]$ of a chart corresponds to a surface point $\mathbf{p}$ of $\mathcal{M}$. Therefore, our goal is to assign a color for each chart pixel $\mathcal{T}[u,v]$. Given a set of multi-view posed images $\mathbf{I}=\{I_{k}\}_{k=1}^{K}$, $\mathbf{p}$ can be seen in zero, one, or more images. If $\mathbf{p}$ is visible in $I_{k}$, we denote $I_{k}$’s corresponding pixel as $I_{k}[i_{k},j_{k}]$. See the top of Figure 4 as an example, where a 3D point $\mathbf{p}$ is visible in two views $a$ and $b$, and we mark it with a red circle together with its corresponding multi-view image pixels $I_{a}[i_{a},j_{a}]$, $I_{b}[i_{b},j_{b}]$, and texture atlas pixel $\mathcal{T}[u,v]$.
Point $\mathbf{p}$’s visibility in view $k$ can calculated by whether its depth value is no bigger than the corresponding pixel of the $k$’s depth map. If $\mathbf{p}$ is visible only in one view, the corresponding color of $I_{k}[i_{k},j_{k}]$ can be directly adopted as the color of $\mathbf{p}$ and $\mathcal{T}[u,v]$. If $\mathbf{p}$ can be seen in no view at all, we can set some rules to deal with it. If $\mathbf{p}$ can be seen in more than one view, we can try to fuse the corresponding colors or select one best view. Various unprojection strategies differ in how to fuse views or select the best view. Here we introduce three implementations for unprojection.
3.3.1. Naive Unprojection: Only Use Direction Priority
@@
While this naive unprojection strategy serves adequately for most points, it produces artifacts under certain circumstances; see again the middle left part of Figure 4, where a gray artifact is visible in the reconstructed textured mesh that should have been brown. Through analysis, such artifacts appear in border areas, that is, the area near the border of occluded (e.g. bottom side of the brown seat in Figure 4) and occluding part (e.g. the gray bottom of the chair) of the shape in a posed image $I_{k}$ when the former is occluded by the latter from view $k$ (e.g. view $a$). In such areas, the inpainting module may inaccurately delineate the boundary separating the two parts, e.g. too large gray bottom in view $a$, resulting in artifacts.
3.3.2. Optimization-based Unprojection
Instead of selecting one single view by direction priority, 3D generation method DreamGaussian (Tang et al., 2023a) adopts an optimization-based unprojection strategy to get the textured mesh from multi-view images. The idea is to optimize the texture atlas by the multi-view per-pixel loss between the inpainted multi-view images $\{I_{k}\}_{k=1}^{K}$ and the rendered images of the reconstructed textured mesh. In this way, the optimization process kinds of fuses corresponding colors in different views instead of selecting the best one. However, it cannot ignore the inaccurate colors in the border areas, so the artifacts can only be reduced instead of eliminated. This inspires us to design a non-border-first strategy to address this issue.
3.3.3. Non-Border-First Unprojection (NBF)
The bottom of Figure 4 illustrates the idea of our proposed novel unprojection strategy namely Non-Border-First Unprojection, which begins with detecting border areas in multi-view images, so as to paint the texture atlas while ignoring them. After that, if some atlas pixels remain unpainted since their associated 3D points can not be seen from any non-border area, we then select from all areas, whether they are border areas or not, to paint the rest pixels. Now the question is, how to accurately detect border areas.
@@ 3.3.3. Non-Border-First Unprojection (NBF)
  3. (3)
    We paint the final texture atlas $\mathcal{T}$ considering only non-border areas in ${\mathcal{T}_{k}}$. During this, if a point is visible in more than one view’s non-border-areas, we select the best view with the highest direction priority.
  4. (4)
    If there remain some unpainted pixels in $\mathcal{T}$, we now check the previously ignored border areas to paint them, during which we still select among views by direction priority.
  5. (5)
@@
    We experimentally find that additionally optimizing from the generated texture atlas by only calculating the loss of non-border pixels would further enhance the performance.
@@
In this way, our non-border-first strategy can reasonably eliminate artifacts caused by inaccurate color prediction in border areas. We conduct experiments to verify the effectiveness of our proposed strategy in Section 4.
4. Experiments and Results
4.1. Experimental Settings
4.1.1. Datasets
@@
To evaluate the performance of our method against other competitors, we use three benchmark datasets in our experiments.
@@
(i) ShapeNetCoreV2 (Chang et al., 2015): A large-scale synthesis 3D object dataset, where we follow (Gao et al., 2022) to use the official test splits of chair, car, and motorbike categories for evaluation since they contain relatively complex textures. The test sets of the three categories contain around 1300, 690 and 70 samples, respectively.
@@
(ii) Google Scanned Objects (GSO) (Downs et al., 2022): A real-scanned 3D object dataset and we use all its 1030 samples for evaluation unlike existing works (Liu et al., 2024; Long et al., 2023) that only use 30 of them.
@@
(iii) OmniObject3D (Wu et al., 2023): A real-scanned 3D object dataset with 6000 samples. For efficiency, we randomly select 100 objects for evaluation.
@@
For each textured mesh from the above datasets, we sample 30000 colored points as input. Note that, our approach works in a zero-shot manner, which can be directly applied to objects from various datasets or categories, requiring no extra training but only an off-the-shelf 2D diffusion model (Dhariwal and Nichol, 2021).
4.1.2. Metrics
@@
To evaluate the quality of the reconstructed 3D textured mesh, we follow existing works (Liu et al., 2023c, 2024; Long et al., 2023; Zhang et al., 2022) to compare the similarity between multi-view images rendered by the reconstructed mesh and ground-truth mesh. Specifically, we render 20 multi-view images on each mesh for a thorough assessment and four similarity metrics are calculated: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) (Wang et al., 2004), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), and Fréchet Inception Distance (FID)  (Heusel et al., 2017). PSNR measures pixel-level similarity. SSIM considers both pixel values and structural information. LPIPS uses a pre-trained deep neural network to assess image similarity, imitating human perception better than PSNR and SSIM. FID measures the statistical distance between the distribution of two datasets. PSNR and SSIM scores indicate higher quality as their values increase, whereas lower values of LPIPS and FID reflect better performance.
4.2. Baselines
@@
To fully evaluate the performance of our method, we compare it with three kinds of baselines:
@@
(1) Non-deep-learning method: We choose the classical Screened Poisson Reconstruction (SPR) (Kazhdan and Hoppe, 2013) for comparison. Note that SPR requires per-point normal as input, and we estimate it via Principal Component Analysis (PCA).
(2) DHSP3D (Wei et al., 2021) and NKSR (Huang et al., 2023): These are two recent networks for textured mesh reconstruction and we employ the official code provided by the authors. Similar to SPR, NKSR also requires point normals as additional input.
@@
(3) Texture Field (Oechsle et al., 2019): Since we cannot find any existing open-source texture-field-based method requiring only colored point clouds as input, we thus design a baseline network inspired by (Gupta et al., 2023). The key idea is to learn a 3D feature tri-plane from the input point cloud and then decode RGB values for the query points. We train it on the official training split of the mentioned three categories of ShapeNetCoreV2, supervised by per-point color MSE loss following (Gupta et al., 2023). During inference, we employ the above-mentioned POCO (Boulch and Marlet, 2022) to reconstruct untextured meshes and utilize our trained texture field network to infer textures by querying colors for the associated 3D point of each texture atlas pixel. In this way, both our proposed PointDreamer and the texture field baseline adopt the same geometry module for a fair comparison. Please refer to our supplementary file for more details about the training and inference of our texture field baseline.
Figure 5. Qualitative results of textured mesh reconstruction on ShapeNetCoreV2 dataset.
Figure 6. Qualitative results of textured mesh reconstruction on Google Scanned Objects dataset.
Figure 7. Qualitative results of textured mesh reconstruction on OmniObject3D dataset.
[…]
Figure 8. Qualitative results on textured mesh reconstruction comparing our method and DHSP3D.
Table 1. Comparisons on textured mesh reconstruction performance on ShapeNetCoreV2 dataset.
@@
Chairs SPR 24.10 0.931 0.0923 12.03
NKSR 23.05 0.931 0.0964 29.11
T.F. 26.12 0.947 0.0664 6.83
Ours 26.29 0.952 0.0570 4.90
Cars SPR
23.46 0.929 0.0882 43.89
NKSR 21.01 0.919 0.0961 131.05
T.F. 22.42 0.919 0.0868 38.08
Ours 22.78
0.930 0.0727 12.41
Motorbikes SPR 15.09 0.809 0.2139 106.09
NKSR 18.13 0.905 0.1008 175.42
T.F. 21.17 0.926 0.0635 40.06
Ours
21.20 0.930 0.0562 28.42
Table 2. Comparisons on textured mesh reconstruction performance on GSO and Omniobject3D datasets. “T.F” denotes Texture Field.

### v2

@@ 3.2 Projection: 3D to 2D
Directly projecting the whole point cloud into 2D space would incorporate points from occluded parts that should not be visible from the given view. This would mess up the inpainting results and subsequent mesh textures; see Figure 2 for an example. Therefore, before generating the sparse images, we process the input point cloud by the “Hidden Point Removal” operator [41], which works by transforming the input and extracting the points that reside on its convex hull. Besides, we compare the depth values of points and the extracted untextured mesh to further remove some invisible points. With pre-set camera parameters of $K$ viewpoints ($K=8$ in our implementation) and associated visible point clouds, we conduct camera transformation to these points to get their corresponding pixel coordinates in 2D image space. These pixels are painted according to their associated 3D points’ colors, producing $K$ sparse images. We compare different $K$ values in our supplementary file.
Algorithm 1 PyTorch-like pseudocode for Non-Border-First (NBF) strategy
1# K: number of views
2# T: texture atlas
3# I: multiview images
4# I_dep: [K, I_res, I_res], multi-view depth maps obtained from geometry
5# T_dep: [K, T_res, T_res], T’s pixels’ corresponding depths in each view
6# Tuv2Iab: [T_res, T_res, 2]: Mapping T’s pixel coordinate to I’s
7# chart_mask: [T_res, T_res]: Does T(u,v) belong to a chart (in foreground)?
8# T_normal: [T_res, T_res, 3]: T’s pixels’ corresponding 3D point’ normal directions
9# view_dir: [K, 3]: camera view directions
10# I_rgb: [K, I_res, I_res, 3], multi-view RGB images
11T = torch.zeros(T_res, T_res, 3)
12vis_T = calc_vis(I_dep, T_dep, Tuv2Iab) # [K, T_res, T_res], bool
13vis_T_edges = detect_edge(vis_T) # [K, T_res, T_res], bool
14chart_edges = detect_edge(chart_mask) # [T_res, T_res], bool
15border_edges = vis_T_edges - chart_edges.unsqueeze(0).repeat(K)
16border_areas = dilate(border_edges, kernel) # [K, T_res, T_res], bool
17cand_view_T_map = non_border_areas = ~border_areas # [K, T_res, T_res]
18Tuv_cand_view_num = non_border_areas.sum(0) # [T_res, T_res], int
19cand_view_T_map[:, Tuv_cand_view_num==0] = vis_T[:, Tuv_cand_view_num==0]
20dir_priority = calc_priority(T_normal, view_dir) # [K, T_res, T_res]
21dir_priority[~cand_view_T_map] -= 1000 # Non-Border-First
22uv_view_id = dir_priority.argmax(0) # [T_res, T_res]
23for u in range(0, T_res):
24 for v in range(0, T_res):
25 T[u, v] = I_rgb[uv_view_id[u, v], Tuv2Iab[u, v, 0], Tuv2Iab[u, v, 1]]
3.3 Inpainting: Sparse to Dense
@@
Any 2D inpainting method that fills empty pixels in input images can serve as our inpainting module. We propose to use the state-or-the-art DDNM [^2] https://github.com/wyhuai/DDNM  [24] based on a pre-trained unconditional 2D diffusion model [^3] https://github.com/openai/guided-diffusion  [38] to inpaint the multi-view sparse images into dense ones $\mathbf{I}=\{I_{k}\}_{k=1}^{K}$. In this way, the 2D diffusion model’s strong prior can facilitate high-quality inpainting. So far, we have completed color prediction purely in 2D space instead of 3D or UV space.
3.4 Unprojection: 2D to 3D
@@
The top of Figure 3 illustrates the basic idea of the unprojection module. Specifically, we represent mesh texture as a texture atlas $\mathcal{T}$ by applying UV mapping to a mesh $\mathcal{M}$. We follow existing works [42, 32] to conduct UV mapping by Xatlas [43]. A texture atlas is an RGB image containing different charts. Each chart is a continuous area in UV space that corresponds to a continuous segment of $\mathcal{M}$ in 3D space. Each pixel $\mathcal{T}[u,v]$ of a chart corresponds to a surface point $\mathbf{p}$ of $\mathcal{M}$. Therefore, our goal is to assign a color for each chart pixel $\mathcal{T}[u,v]$. Given a set of multi-view posed images $\mathbf{I}=\{I_{k}\}_{k=1}^{K}$, $\mathbf{p}$ can be seen in zero, one, or more images. If $\mathbf{p}$ is visible in $I_{k}$, we denote $I_{k}$’s corresponding pixel as $I_{k}[i_{k},j_{k}]$. See the top of Figure 3 as an example, where a 3D point $\mathbf{p}$ is visible in two views $a$ and $b$, and we mark it with a red circle together with its corresponding multi-view image pixels $I_{a}[i_{a},j_{a}]$, $I_{b}[i_{b},j_{b}]$, and texture atlas pixel $\mathcal{T}[u,v]$.
Point $\mathbf{p}$’s visibility in view $k$ can be calculated by whether its depth value is no bigger than the corresponding pixel of the $k$’s depth map. If $\mathbf{p}$ is visible only in one view, the corresponding color of $I_{k}[i_{k},j_{k}]$ can be directly adopted as the color of $\mathbf{p}$ and $\mathcal{T}[u,v]$. If $\mathbf{p}$ can be seen in no view at all, we can set some rules to deal with it. If $\mathbf{p}$ can be seen in more than one view, we can try to fuse the corresponding colors or select one best view. Various unprojection strategies differ in how to fuse views or select the best view. Here we introduce three different implementations for unprojection.
3.4.1 Naive Unprojection: Only Use Direction Priority
@@
While this naive unprojection strategy serves adequately for most points, it produces artifacts under certain circumstances; see again the middle left part of Figure 3, where a gray artifact is visible in the reconstructed textured mesh that should have been brown. Through analysis, such artifacts appear in border areas: the area near the border of occluded and occluding part of the shape when viewing from view $k$. For example, in view $a$, the brown seat’s bottom side is occluded by the gray bottom of the chair. The area near the edge of the brown and gray parts is a border area. In such areas, the inpainting module and the geometry extraction module may inconsistently delineate the boundary separating the two parts. For example, the gray bottom is larger in the inpainted image $I_{a}$ than of the reconstructed mesh, thus resulting in an artifact.
3.4.2 Optimization-based Unprojection
Instead of selecting one single view by direction priority, 3D generation method DreamGaussian [19] adopts an optimization-based strategy to get the textured mesh from multi-view images. The idea is to optimize the texture atlas by a per-pixel loss between the inpainted multi-view images $\{I_{k}\}_{k=1}^{K}$ and the rendered images of the reconstructed textured mesh. In this way, the optimization process kinds of fuses corresponding colors in different views instead of selecting the best one. However, it cannot ignore the inconsistent colors in the border areas, so the artifacts can only be reduced instead of eliminated. This inspires us to design a non-border-first strategy to address this issue.
3.4.3 Non-Border-First Unprojection (NBF)
The bottom of Figure 3 illustrates the idea of our proposed novel unprojection strategy namely Non-Border-First Unprojection. It begins with detecting border areas in multi-view images, so as to paint the texture atlas while deprioritizing them. After that, if some atlas pixels remain unpainted since their associated 3D points can not be seen from any non-border area, we then select from all areas, whether they are border areas or not, to paint the rest pixels. Now the question is, how to accurately detect border areas.
@@ 3.4.3 Non-Border-First Unprojection (NBF)
  3. 3.
    We paint the final texture atlas $\mathcal{T}$ considering only non-border areas in ${\mathcal{T}_{k}}$. During this, if a point is visible in more than one view’s non-border areas, we select the view with the highest direction priority.
  4. 4.
    If there remain some unpainted pixels in $\mathcal{T}$, check the previously ignored border areas to paint them, during which we still select among views by direction priority.
  5. 5.
@@
    We experimentally find that additionally optimizing from the generated texture atlas by only calculating the loss of non-border pixels would further enhance the results.
@@
In this way, our non-border-first strategy can reasonably eliminate artifacts caused by inconsistent prediction in border areas. Altorithm 1 shows the corresponding pseudoCode. We conduct experiments to verify the effectiveness of our proposed strategy in Section 4.
4 Experiments and Results
4.1 Experimental Settings
4.1.1 Datasets
@@
To evaluate the performance of our method against other competitors, we use three benchmark datasets:
@@
(i) ShapeNetCoreV2 [44]: A large-scale synthesis 3D object dataset, where we follow [42] to use the official test splits of chair, car, and motorbike categories for evaluation, since they contain relatively complex textures. The test sets of the three categories contain around 1300, 690, and 70 samples, respectively.
@@
(ii) Google Scanned Objects (GSO) [45]: A real-scanned 3D object dataset and we use all its 1030 samples for evaluation, unlike existing works [14, 32] that only use 30 of them.
@@
(iii) OmniObject3D [46]: A real-scanned 3D object dataset with 6000 samples. For efficiency, we randomly select 100 objects for evaluation.
Figure 4: Qualitative results on (a) ShapeNetCoreV2 dataset (Chair, Car, Motorbike Category), (b) GSO dataset, (c) OmniObject3D dataset. (d) Comparison with DHSP on GSO. (e) Our Multiview Consistency.
@@
For each textured mesh from the above datasets, we sample 30k colored points as input. Note that, our approach is zero-shot and can be directly applied to objects from various datasets or categories, requiring no extra training but only an off-the-shelf 2D diffusion model [38].
4.1.2 Metrics
@@
To evaluate the quality of the reconstructed 3D textured mesh, we follow existing works [22, 14, 32, 47] to compare the similarity between multi-view images rendered by the reconstructed mesh and ground-truth mesh. Specifically, we render 20 images per shape for a thorough assessment by distributing cameras on the vertices of a regular icosahedron. Four similarity metrics are calculated: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [48], Learned Perceptual Image Patch Similarity (LPIPS) [49], and Fréchet Inception Distance (FID)  [50]. PSNR and SSIM focus more on pixel-level similarity, while LPIPS and FID imitate human perception better. Since we focus on texture instead of geometry reconstruction, we put geometry evaluation results in our supplementary file.
4.2 Baselines
@@
To fully evaluate the performance of our method, we compare it with three kinds of baseline methods:
@@
(1) Screened Poisson Reconstruction (SPR) [7]: SPR is a classical mesh reconstruction method. We adopt the implementation of the commonly-used tool MeshLab [26], which generates texture by linearly blending colors of the input point cloud. Note that SPR requires per-point normal as input, which we estimate via MeshLab’s default Principal Component Analysis (PCA).
(2) DHSP3D [8] and NKSR [9]: These are two recent approaches for mesh reconstruction which generate textures by overfitting the input points’ colors. We employ the official code provided by the authors  [^4] https://github.com/weixk2015/DHSP3D,
https://github.com/nv-tlabs/NKSR/blob/public/examples/
recons_colored_mesh.py
. Similar to SPR, NKSR also requires point normals as additional input.
@@
(3) Texture Field (T.F.)  [12]: Since we cannot find any existing open-source texture-field-based method requiring only colored point clouds as input, we thus design a baseline network inspired by [10]. The key idea is to learn a 3D feature tri-plane from the input point cloud and then decode RGB values for the query points. We train it on the official training split of the mentioned three categories of ShapeNetCoreV2, supervised by per-point color MSE loss following [10]. During inference, we employ the above-mentioned POCO [4] to reconstruct untextured meshes. Then we utilize our trained T.F. network to infer textures, by querying colors for the associated 3D point of each texture atlas pixel. In this way, both PointDreamer and the T.F. baseline adopt the same geometry module for a fair comparison. Please refer to our supplementary file for more details about the training and inference of our T.F. baseline.
TABLE I: Comparisons on textured mesh reconstruction on ShapeNetCoreV2, GSO and Omniobject3D datasets. T.F. refers to Texture Field. Note that both T.F. and our PointDreamer adopt POCO [4] for geometry reconstruction for a fair comparison.
@@
Chairs SPR 24.10 0.931 0.092 12.03
NKSR 23.05 0.931 0.096 29.11
T.F. 26.12 0.947 0.066 6.83
Ours 26.29 0.952 0.057 4.90
Cars SPR 23.46 0.929 0.088 43.89
NKSR 21.01 0.919 0.096 131.05
T.F. 22.42 0.919 0.087 38.08
Ours 22.78 0.930 0.073 12.41
Motorbikes SPR 15.09 0.809 0.214 106.09
NKSR 18.13 0.905 0.101 175.42
T.F. 21.17 0.926 0.064 40.06
Ours 21.20 0.930 0.056 28.42
Dataset Method PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
GSO SPR 27.11 0.907 0.120 27.43
NKSR 24.65 0.900 0.118 36.51
T.F. 25.53 0.894 0.128 46.44
Ours 27.24 0.923 0.083 9.32
OmniObject3D SPR 30.20 0.927 0.100 40.13
NKSR 26.77 0.919 0.105 50.14
T.F. 28.08 0.914 0.112 65.61
Ours 29.78 0.941 0.068 18.23
4.3 Main Results
4.3.1 Qualitative Results
We present visual results in Figure 4. Since DHSP3D is much slower than other methods (4-6 hours per shape compared to $<$1 min), we randomly select two objects from GSO dataset for efficiency during comparison. We highly recommend referring to our supplementary materials for more visual results and a 360° video.
Figure 4 (a)-(d) show that our method, though zero-shot, achieves much clearer and more realistic textures than baselines, thanks to the 2D diffusion prior. Besides, SPR and NKSR, which both rely on per-point normals as input, sometimes yield redundant geometry mainly due to the wrongly-estimated normal direction. The texture field baseline, which is trained on ShapeNetCoreV2, performs much worse on the newly-seen dataset of GSO and Omniobject3D, indicating a lack of generalization ability. In Figure 4 (d), the red cup in the first row shows that DHSP3D does not support arbitrary topology. This is because it optimizes meshCNN and 2D XYZ map from the convex hull of the input point cloud, enforcing the output mesh to adhere to the hull’s topology. Besides, DHSP3D suffers from less realistic textures with jagged or unclear edges, since its purely self-prior property only enables it to predict point colors considering corresponding neighbors, without the ability to dream the unseen like our PointDreamer.
4.3.2 Quantitative comparisons
We summarize the quantitative results in Tables 4.2, and geometry evaluation results in our supplementary file. Our method outperforms baselines on most metrics and datasets, especially on the two perceptual metrics FID and LPIPS with a significant margin. However, regarding PSNR, which prioritizes pixel-level accuracy and thus may differ from human perception of quality, our method is sometimes outperformed by SPR. PSNR has been found to prefer blurry images [49]. We assume that SPR predicts a point’s color by blending nearby points’ colors, leading to blurry textures but avoiding extreme pixel-level errors. On the contrary, our method dreams the colors by its diffusion prior. While achieving visually better results, a few pixels with extremely high errors may lower the PSNR score.
4.4 Multi-view Consistency
Figure 4 (e) and our supplementary video show that, our method achieves high multi-view consistency. The reason is two-fold: (1) Unlike image-conditioned 3D generation, it’s easier to maintain multi-view consistency when reconstructing point clouds, since the input point cloud already determines the global shape and colors, leaving only local details to be dreamed by the model. (2) For inconsistencies in local details, our proposed Non-Border-First unprojection effectively addresses this issue; see Figure 7 (c) and Table 4.4.
TABLE II: Anti-noise comparisons on ShapeNetCoreV2 Chairs. ‘Noisy’ means adding Gaussian noise with a standard deviation of 0.005 to the input.